OpenStudy (anonymous):

A recipe requires 13 oz. of sugar and 18 oz. of flour. If only 10 oz. of sugar are used, how much flour, to the nearest ounce, should be used?

OpenStudy (anonymous):

13 oz. is the typical amount of sugar 18 oz. is the typical amount of flour 10 oz. is the amount of sugar used Let x = the amount of flour that should be used for 10 oz. of sugar This question is probably a proportion question, so basically what you would so is put the typical amount of sugar (13 oz.) divided by the typical amount of flour (18 oz.), and make that equal to the amount of sugar used (10 oz.) divided by the amount of flour you should use (x). 13/18 = 10/x Then, use cross products to solve for x. 13x = 18*10 13x=180 x=180/13 or approximately 13.8 oz. of flour.